Greydolphin15:

Jose asks his friends to guess the higher of two grades he received on his math tests. He gives them two hints. The difference between the two grades is 16. The sum of one-eighth of the higher grade and one-half of the lower grade is 52. The system that represents his scores is below. x – y = 16 1/8x+1/2y=52 What is the higher grade of Jose’s two tests? 48 52 80 96

darkknight:

x-y=16 x = y+16

darkknight:

1/8(y+16) + 1/2(y) = 52

darkknight:

1/8 y + 2 +1/2 y = 52 5/8 y = 50 y = 50(8/5) y = 80

Greydolphin15:

Thank youuu

darkknight:

plug back in and solve for x

darkknight:

x is the answer because 80+16 = x which is 96

darkknight:

np